OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (1 point) 2x + 9y = 4 3x + 7y = –7 Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (1 point) 5x – y = –2 10x – 3y = –7

OpenStudy (texaschic101):

I would multiply the 1st equation by -3 and the 2nd equation by 2. By doing this, the x's will cancel out (thus the elimination method). 2x + 9y = 4 -->(-3) 3x + 7y = -7 -->(2) ------------------ -6x - 27y = -12 (result of multiplying by -3) 6x + 14y = -14 (result of multiplying by 2) ------------------add 0 - 13y = - 26 -13y = -26 y = -26/-13 (negative divided by negative equals positive) y = 2 Now sub 2 in for y in either of the original equations 2x + 9y = 4 2x + 9(2) = 4 2x + 18 = 4 2x = -18 + 4 2x = - 14 x = -14/2 x = -7 you can now check your answers 3x + 7y = -7 2x + 9y = 4 3(-7) + 7(2) = -7 2(-7) + 9(2) = 4 -21 + 14 = -7 -14 + 18 = 4 -7 = -7 (correct) 4 = 4 (correct) ====================== You want to pick an equation and isolate a variable in it. Then sub what that variable equals into the other equation and solve for the other variable. 5x - y = -2 -y = -2 - 5x (multiply by -1 to make y positive) y = 5x + 2 now sub 5x + 2 in for y in the other equation 10x - 3y = -7 10x - 3(5x + 2) = -7 10x - 15x - 6 = -7 -5x - 6 = -7 -5x = -7 + 6 -5x = -1 x = -1/-5 x = 1/5 now sub 1/5 in for x in the 1st equation 5x - y = -2 5(1/5) - y = -2 1 - y = -2 -y = -2-1 -y = -3 (multiply by -1 to make y positive) y = 3 now check your answers 5x - y = -2 10x - 3y = -7 5(1/5) - 3 = -2 10(1/5) - 3(3) = -7 1 - 3 = -2 2 - 9 = -7 -2 = -2 (correct) -7 = -7 (correct)